Prove with Boolean algebra that (x - y) + (x-y)-y. Give a reason for each step in your proof.
Answer #1

= x.y+x'.y

= y(x+x')

= y(1) [Since x+x' = 1]

= y

Therefore x.y+x'.y = y

____________Thank You